OpenStudy (anonymous):

You are offered two different sales jobs. The first company offers a straight commission of 8% of the sales. The second company offers a salary of $ 380 per week plus 4% of the sales. How much would you have to sell in a week in order for the straight commission offer to be at least as good? The answer to this question was 9500, meaning you received $760 in commission of the $9500 worth of sales for the week. I do not understand this answer. How do we know how many sales were made? I was doing the problem assuming 0 sales were made since it didn't specify. Brain fart! Can someone point me

OpenStudy (jdoe0001):

if you were to write ... say 8% in decimal.... how would it look like?

OpenStudy (jdoe0001):

keep in mind that \(\bf {\color{red}{ a}}\%\ of \ b=\cfrac{{\color{red}{ a}}}{100}\times b\)

OpenStudy (anonymous):

yeah, 8% is obviously .08, this is why I am confused. I would think that the answer would be 4750, since if you were to make 4750 in sales, you would receive $380 in commission. However, the answer is $9500, and I do not understand why.

OpenStudy (anonymous):

Let s=sales 1st company salary = 8% * s = .08s 2nd company salary = 380 + 4% * s = 380+ .04s Since the 1st company salary and 2nd company salary should be equal, we get the equation: .08s = 380 + .04s Can you solve that? Also sorry for taking so long to type, my computer was lagging really badly

OpenStudy (anonymous):

No worries! ^ thank you, that helps a lot

OpenStudy (anonymous):

Np :)